in 1980, the median family income was 18,000 and in 2000 it was 41,000 I need to know what the m= and what would be the median family income 2002
Answer by stanbon(75887)

in 1980, the median family income was 18,000 and in 2000 it was 41,000 I need to know what the m= and what would be the median family income 2002
----
Let "x" be the number of years after 1980.
----
You have 2 points: (0,18000) and (20,41000)
----
Find an equation thru those two points:
slope = (41000-18000)/(20-0) = 23000/20 = 1150
---
Intercept = f(0) = 18000
-------
Equation:
f(x) = 1150x + 18000
----
Income in 2002
--
f(22) = 1150*22 + 18000
f(22) = $43300
=====================
Cheers,
Stan H.